# Deep FlexQP: Accelerated Nonlinear Programming via Deep Unfolding

### v1

@@ Abstract
We propose an always-feasible quadratic programming (QP) optimizer, FlexQP, which is based on an exact relaxation of the QP constraints. If the original constraints are feasible, then the optimizer finds the optimal solution to the original QP. On the other hand, if the constraints are infeasible, the optimizer identifies a solution that minimizes the constraint violation in a sparse manner. FlexQP scales favorably with respect to the problem dimension, is robust to both feasible and infeasible QPs with minimal assumptions on the problem data, and can be effectively warm-started. We subsequently apply deep unfolding to improve our optimizer through data-driven techniques, leading to an accelerated Deep FlexQP. By learning dimension-agnostic feedback policies for the parameters from a small number of training examples, Deep FlexQP generalizes to problems with larger dimensions and can optimize for many more iterations than it was initially trained for. Our approach outperforms two recently proposed state-of-the-art accelerated QP approaches on a suite of benchmark systems including portfolio optimization, classification, and regression problems. We provide guarantees on the expected performance of our deep QP optimizer through probably approximately correct (PAC) Bayes generalization bounds. These certificates are used to design an accelerated sequential quadratic programming solver that solves nonlinear optimal control and predictive safety filter problems faster than traditional approaches. Overall, our approach is very robust and greatly outperforms existing non-learning and learning-based optimizers in terms of both runtime and convergence to the optimal solution across multiple classes of NLPs.
*Work done while at Georgia Tech.
@@ 1 Introduction
Nonlinear programming (NLP) is a key technique for both large-scale decision making, where difficulty arises due to the sheer number of variables and constraints, as well as real-time embedded systems, which need to solve many NLPs with similar structure quickly and robustly. Within NLP, quadratic programming (QP) plays a fundamental role as many real-world problems in optimal control (anderson2007optimal), portfolio optimization (markowitz1952portfolio; boyd2013performance; boyd2017multi), and machine learning (huber1964robust; cortes1995support; tibshirani1996regression; candes2008enhancing) can be represented as QPs.
Furthermore, sequential quadratic programming (SQP) methods utilize QP as a submodule to solve much more complicated problems where the objective and constraints may be nonlinear and non-convex, such as in nonlinear model predictive control (diehl2009efficient; rawlings2020model), state estimation (aravkin2017generalized), and power grid optimization (montoya2019sequential). SQP itself can even be used as a subproblem for solving mixed integer NLPs (leyffer2001integrating) and large-scale partial differential equations (fang2023sqp).
@@
However, a common difficulty with SQP methods occurs when the linearization of the constraints results in an infeasible QP subproblem, and a large amount of research has focused on how to repair or avoid these infeasibilities, e.g., (fletcher1985l1; izmailov2012stabilized), among others. A significant advantage of SNOPT (gill2005snopt), one of the most well-known SQP-based methods, is in its infeasibility detection and reduction handling. These considerations necessitate a fast yet robust QP solver that works under minimal assumptions on the problem parameters.
@@
To this end, we propose FlexQP, a flexible QP solver that is always-feasible, meaning that it can solve any QP regardless of the feasibility of the constraints. Our method is based on an exact relaxation of the QP constraints: if the original QP was feasible, then FlexQP will identify the optimal solution. On the other hand, if the original QP was infeasible, instead of erroring or failing to return a solution, FlexQP automatically identifies the infeasibilities while simultaneously finding a point that minimizes the constraint violation. This allows FlexQP to be a robust QP solver in and of itself, but its power shines when used a submodule in an SQP-type method, see Figure 1.
@@
Figure 1: SQP with Deep FlexQP can solve highly-constrained nonlinear optimizations over 15x faster than SQP with OSQP (averaged over 100 problems).
@@
Moreover, through the relaxation of the constraints, multiple hyperparameters are introduced that can be difficult to tune and have a non-intuitive effect on the optimization. To address this shortcoming, we use deep unfolding (monga2021algorithm) to design lightweight feedback policies for the parameters based on actual problem data and solutions for QP problems of interest, leading to an accelerated version titled Deep FlexQP. Learning the parameters in a data-driven fashion avoids the laborious process of tuning them by hand or designing heuristics for how they should be updated from one iteration to the next. Meanwhile, these data-driven rules have been shown to strongly outperform the hand-crafted ones, such as in the works by ichnowski2021accelerating and saravanos2025deep.
@@
We thoroughly benchmark Deep FlexQP against traditional and learned QP optimizers on multiple QP problem classes including machine learning, portfolio optimization, and optimal control problems. Moreover, we certify the performance of Deep FlexQP through probably approximately correct (PAC) Bayes generalization bounds, which provide a guarantee on the mean performance of the optimizer. We propose a log-scaled training loss that better captures the performance of the optimizer when the residuals are very small. Finally, we deploy Deep FlexQP to solve nonlinearly-constrained trajectory optimization and predictive safety filter problems (wabersich2021predictive). Overall, Deep FlexQP can produce an order-of-magnitude speedup over OSQP (stellato2020osqp) when deployed as a subroutine in an SQP-based approach (Figure 1), while also robustly handling infeasibilities that may occur due to a poor linearization or an over-constrained problem.
@@ 2 Motivation & Related Work
$\begin{aligned} \operatorname*{minimize}_{x}\hskip 5.0pt&f(x),\\ \operatorname*{subject\ to}\hskip 5.0pt&g(x)\leq 0,\\ &h(x)=0,\end{aligned}$ (1)
@@
where $f:\mathbb{R}^{n}\to\mathbb{R}$ twice-differentiable is the objective to be minimized and $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{p}$ differentiable describe the inequality and equality constraints, respectively. SQP solves Equation 1 by iteratively linearizing the constraints and quadraticizing the Lagrangian $\mathcal{L}(x,y_{I},y_{E})\vcentcolon=f(x)+y_{I}^{\top}g(x)+y_{E}^{\top}h(x)$ around the current iterate $(x^{k},y_{I}^{k},y_{E}^{k})$, where $y_{I}\in\mathbb{R}^{m}_{+}$ and $y_{E}\in\mathbb{R}^{p}$ are the dual variables for the inequality and equality constraints, respectively. This results in the following QP subproblem:
$\operatorname*{minimize}_{dx}\hskip 5.0pt$ $\frac{1}{2}dx^{\top}\nabla_{x}^{2}\mathcal{L}(x^{k},y_{I}^{k},y_{E}^{k})\,dx+\nabla f(x^{k})^{\top}dx,$ (2a)
$\operatorname*{subject\ to}\hskip 5.0pt$ $g(x^{k})+\partial g(x^{k})dx\leq 0,$ (2b)
$h(x^{k})+\partial h(x^{k})dx=0.$ (2c)
@@
Notably, the linearization of the constraints $g$ and $h$ may not produce a QP subproblem that is feasible, meaning that there may not exist any $dx$ that satisfies the linearized constraints Equations 2b and 2c. In this case, the SQP solver either terminates with a suboptimal point or a specialized routine needs to be run in order to reduce the infeasibility. For example, when SNOPT encounters an infeasible subproblem, it enters elastic mode and solves a new optimization where the constraints are relaxed using $\ell_{1}$ penalty functions (gill2005snopt). This is advantageous over other choices, such as an $\ell_{2}$ penalty, as the $\ell_{1}$ norm encourages sparsity in the constraint violation. This means the optimizer can naturally identify the constraints that are the most difficult to satisfy. Similarly, stabilized SQP methods attempt to regularize the constraints at every step, but this requires determining a suitable regularization through another process (wright1998superlinear; hager1999stabilized). In the context of mixed integer NLPs, infeasibilities are very likely to occur during the branch and bound process, so their fast and robust identification is crucial (gill2011sequential).
@@
Moreover, as the interest in data-driven optimization grows, we often wish to solve many optimization problems with similar structure repeatedly (amos2023tutorial), and potentially in parallel or in a batched fashion, such as in the method by fang2023sqp, which requires solving coupled systems of SQP problems in parallel. Needing to run specialized procedures in these cases is not scalable. Furthermore, identifying the best hyperparameters for each individual optimization problem is difficult and time consuming. Deep unfolding is a learning-to-optimize approach (chen2022learning) that has roots in the signal and image processing domains (gregor2010ista; wang2015deep) and utilizes data-driven machine learning to reduce the cost of hyperparameter tuning and to accelerate convergence of model-based optimizers (monga2021algorithm). It constitutes the state-of-the-art approach for sparse recovery (liu2019alista) and video reconstruction (deweerdt2024transformer). In the context of NLP, deep unfolding has been recently applied to accelerate QPs. saravanos2025deep use an analogy to closed-loop control and learn feedback policies for the parameters of a deep-unfolded variant of the operator splitting QP (OSQP) solver (stellato2020osqp), which is a first-order method based on the alternating direction method of multipliers (ADMM) (boyd2011distributed). Their method can achieve orders-of-magnitude improvement in wall-clock time compared to OSQP, and they also propose a decentralized version for quickly solving QPs with distributed structure. Their idea is similar in vein to that of ichnowski2021accelerating, who use reinforcement learning to train a policy that outputs the optimal parameters for OSQP, with the goals of accelerating the optimizer. Another related approach learns to warm-start a Douglas-Rachford splitting QP solver, with the goal of improving convergence speed (sambharya2023end).
Based on these considerations, in Section 3 we propose a QP optimizer, FlexQP, that robustly handles infeasible QPs in a unified manner. We subsequently apply deep unfolding in Section 4, leading to a data-accelerated variant, Deep FlexQP, that outperforms traditional optimization on multiple classes of NLP problems (Figures 1 and 5).
@@ 3 FlexQP: An Always-Feasible Quadratic Programming Solver
Our proposed QP solver, FlexQP, transforms the QP constraints using an exact relaxation and then solves the resultant problem using an operator splitting inspired by OSQP (stellato2020osqp). We will assume the reader is familiar with ADMM; a good overview is provided by boyd2011distributed.
@@ 3.1 Quadratic Programming
We are interested in solving QPs of the general form
$\operatorname*{minimize}_{x}\hskip 5.0pt$ $\frac{1}{2}x^{\top}Px+q^{\top}x,$ (3a)
$\operatorname*{subject\ to}\hskip 5.0pt$ $Gx\leq h,$ (3b)
$Ax=b,$ (3c)
@@
The optimality conditions for Equation 3 are given by:
$Px+q+G^{\top}y_{I}+A^{\top}y_{E}$ $=0,$ (4a)
$Gx-h$ $\leq 0,$ (4b)
$Ax-b$ $=0,$ (4c)
$y_{I}\odot(Gx-h)$ $=0,$ (4d)
$y_{I}$ $\geq 0,$ (4e)
where $y_{I}\in\mathbb{R}^{m}$ and $y_{E}\in\mathbb{R}^{p}$ are the dual variables for the inequality and equality constraints, respectively. A tuple $(x^{*},y_{I}^{*},y_{E}^{*})$ satisfying Equation 4 is a solution to Equation 3.
Throughout this work, we will avoid making any assumptions on $G$ or $A$, meaning that the constraints may be redundant, and in the worst case, there may not exist a feasible $x$ for the optimization. This allows for a unified way to handle infeasibilities when optimizations of the form Equation 3 are embedded as a subproblem in SQP, as in Equation 2.
@@ 3.2 Elastic Formulation
By introducing slack variables $s\in{\mathbb{R}}^{m}$, Equation 3 can be expressed equivalently as
$\operatorname*{minimize}_{x,s\geq 0}\hskip 5.0pt$ $\frac{1}{2}x^{\top}Px+q^{\top}x,$ (5a)
$\operatorname*{subject\ to}\hskip 5.0pt$ $Gx+s-h=0,$ (5b)
$Ax-b=0.$ (5c)
This standard technique is the basis of many interior point algorithms (nocedal2006numerical). We then relax the set of equality constraints Equations 5b and 5c using $\ell_{1}$ penalty functions, yielding
@@
with elastic penalty parameters $\mu_{I},\mu_{E}>0$. This relaxation approach is known as elastic programming (brown1975elastic), and one of the most well-known SQP-based solvers, SNOPT, uses this technique in order to reduce the infeasibility of a QP subproblem (gill2005snopt). This relaxation is also a fundamental step in the sequential $\ell_{1}$ quadratic programming method of fletcher1985l1. Notably, if Equation 3 has a feasible solution and the elastic penalty parameters are sufficiently large, then the solutions to Equation 3 and Equation 6 are identical — this is why the relaxation is exact. On the other hand, if the original QP Equation 3 is infeasible, then solving Equation 6 finds a point that minimizes the constraint violation (nocedal2006numerical). This is formalized through the following theorem, which also describes what we mean by a sufficiently large penalty parameter.
@@
Let $(x^{*},y_{I}^{*},y_{E}^{*})$ solve Equation 3. Let $\mu_{I}^{*}=\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}^{*}=\left\lVert y_{E}^{*}\right\rVert_{\infty}$. Then, for all $\mu_{I}\geq\mu_{I}^{*}$ and $\mu_{E}\geq\mu_{E}^{*}$, the minimizers of Equation 3 and Equation 6 coincide.
This theorem is a generalization of the one by han1979exact that shows we can select a different penalty parameter for the inequality vs. equality constraints. The proof, provided in Appendix A relies on two simple facts: the optimality conditions of Equation 3 and the convexity of the objective. The proof also shows that it is possible to select vectors of penalty parameters $\mu_{I}$ and $\mu_{E}$, as long as each $\mu_{I,i}$ and $\mu_{E,j}$ obeys the constraints $\mu_{I,{i}}\geq|y_{I,i}|$ or $\mu_{E,j}\geq|y_{E,j}|$, respectively.
What happens when the penalty parameters $\bm{\mu}$ do not satisfy the conditions of Theorem 3.1? Using the interpretation of the Lagrange multiplier $y_{i}$ representing the cost of an individual constraint $i$ with associated penalty parameter $\mu_{i}>0$, if $\mu_{i}\geq|y_{i}|$ then the penalty on violating constraint $i$ in Equation 6 is large enough such that Theorem 3.1 holds. On the other hand, if $\mu_{i}<|y_{i}|$, then this constraint $i$ is not being penalized strong enough, and so the solution to Equation 6 will violate this constraint, with the amount of violation proportional to the difference between $\mu_{i}$ and $|y_{i}|$. We use this interpretation in Section 4 to design feedback policies that select the best penalty parameters as a function of the optimizer state and enforce the condition $\mu_{i}\geq|y_{i}|$ during learning using a supervised loss that includes the Lagrange multipliers (see also Theorem 3.3 below).

### v2

@@ Abstract
We propose FlexQP, an always-feasible convex quadratic programming (QP) solver based on an $\ell_{1}$ elastic relaxation of the QP constraints. If the original constraints are feasible, FlexQP provably recovers the optimal solution. If the constraints are infeasible, FlexQP identifies a solution that minimizes the constraint violation while keeping the number of violated constraints sparse. Such infeasibilities arise naturally in sequential quadratic programming (SQP) subproblems due to the linearization of the constraints. We prove the convergence of FlexQP under mild coercivity assumptions, making it robust to both feasible and infeasible QPs. We then apply deep unfolding to learn LSTM-based, dimension-agnostic feedback policies for the algorithm parameters, yielding an accelerated Deep FlexQP. To preserve the exactness guarantees of the relaxation, we propose a normalized training loss that incorporates the Lagrange multipliers. We additionally design a log-scaled loss for PAC-Bayes generalization bounds that yields substantially tighter performance certificates, which we use to construct an accelerated SQP solver with guaranteed QP subproblem performance. Deep FlexQP outperforms state-of-the-art learned QP solvers on a suite of benchmarks including portfolio optimization, classification, and regression problems, and scales to dense QPs with over 10k variables and constraints via fine-tuning. When deployed within SQP, our approach solves nonlinear trajectory optimization problems 4-16x faster than SQP with OSQP while substantially improving success rates. On predictive safety filter problems, Deep FlexQP reduces safety violations by over 70% and increases task completion by 43% compared to existing methods.
@@ 1 Introduction
Nonlinear programming (NLP) is a key technique for both large-scale decision making, where difficulty arises due to the sheer number of variables and constraints, and real-time embedded systems, which need to solve many NLPs with similar structure quickly and robustly. Within NLP, quadratic programming (QP) plays a fundamental role as many real-world problems in optimal control (Anderson and Moore, 2007), portfolio optimization (Markowitz, 1952; Boyd et al., 2013; 2017), and machine learning (Huber, 1964; Cortes and Vapnik, 1995; Tibshirani, 1996; Candes et al., 2008) can be represented as QPs. Furthermore, sequential quadratic programming (SQP) methods utilize QP as a submodule to solve much more complicated problems where the objective and constraints may be nonlinear and nonconvex, such as in nonlinear model predictive control (MPC) (Diehl et al., 2009; Rawlings et al., 2020), state estimation (Aravkin et al., 2017), and power grid optimization (Montoya et al., 2019). SQP itself can even be used as a subproblem for solving mixed integer NLPs (Leyffer, 2001) and large-scale partial differential equations (Fang et al., 2023).
@@
However, a common difficulty with SQP methods occurs when the linearization of the constraints results in an infeasible QP subproblem, and a large amount of research has focused on how to repair or avoid these infeasibilities, e.g., (Fletcher, 1985; Izmailov and Solodov, 2012), among others. A significant advantage of SNOPT (Gill et al., 2005), one of the most well-known SQP-based methods, is in its infeasibility detection and reduction handling. These considerations necessitate a fast yet robust QP solver that works under minimal assumptions on the problem parameters.
@@
To this end, we propose FlexQP, a flexible QP solver that is always-feasible, meaning that it can solve any QP regardless of the feasibility of the constraints. Our method is based on an exact relaxation of the QP constraints: if the original QP is feasible, then FlexQP identifies the optimal solution. On the other hand, if the original QP is infeasible, instead of erroring or failing to return a solution, FlexQP automatically identifies the infeasibilities while simultaneously finding a point that minimizes the constraint violation. This allows FlexQP to be a robust QP solver in and of itself, but its power shines when used as a submodule in an SQP-type method, see Fig. 1.
@@
Figure 1: SQP with Deep FlexQP can solve highly constrained nonlinear optimizations over 16x faster than SQP with OSQP (averaged over 100 problems).
@@
Moreover, through the relaxation of the constraints, multiple hyperparameters are introduced that can be difficult to tune and have a non-intuitive effect on the optimization. To address this shortcoming, we use deep unfolding (Monga et al., 2021) to design lightweight feedback policies for the parameters based on actual problem data and solutions for QP problems of interest, leading to an accelerated version titled Deep FlexQP. Learning the parameters in a data-driven fashion avoids the laborious process of tuning them by hand or designing heuristics for how they should be updated from one iteration to the next. Meanwhile, these data-driven rules have been shown to strongly outperform the hand-crafted ones, such as in the works by Ichnowski et al. (2021) and Saravanos et al. (2025).
@@
We thoroughly benchmark Deep FlexQP against traditional and learned QP optimizers on multiple QP problem classes including machine learning, portfolio optimization, and optimal control problems. Moreover, we certify the performance of Deep FlexQP through probably approximately correct (PAC) Bayes generalization bounds, which provide a guarantee on the mean performance of the optimizer. We propose a log-scaled training loss that better captures the performance of the optimizer when the residuals are very small. Finally, we deploy Deep FlexQP to solve nonlinearly constrained trajectory optimization and predictive safety filter problems (Wabersich and Zeilinger, 2021). Overall, Deep FlexQP can produce an order-of-magnitude speedup over OSQP (Stellato et al., 2020) when deployed as a subroutine in an SQP-based approach (Fig. 1), while also robustly handling infeasibilities that may occur due to a poor linearization or an over-constrained problem.
@@ 2 Motivation & Related Work
$\operatorname*{minimize}_{x}\hskip 5.0ptf(x),\quad\operatorname*{subject\ to}\hskip 5.0ptg(x)\leq 0,\hskip 5.0pth(x)=0,$ (1)
@@
where $f:\mathbb{R}^{n}\to\mathbb{R}$ twice-differentiable is the objective to be minimized and $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{p}$ twice-differentiable describe the inequality and equality constraints, respectively. SQP solves Eq. 1 by solving a sequence of QP problems that are produced by linearizing the constraints and quadraticizing the Lagrangian [^3] In practice, the Hessian of the Lagrangian $\nabla_{x}^{2}\mathcal{L}$ may be indefinite and is often replaced by a positive definite approximation, e.g., via a quasi-Newton update. $\mathcal{L}(x,y_{I},y_{E})\vcentcolon=f(x)+y_{I}^{\top}g(x)+y_{E}^{\top}h(x)$ around the current iterate $(x^{k},y_{I}^{k},y_{E}^{k})$, where $y_{I}\in\mathbb{R}^{m}_{+}$ and $y_{E}\in\mathbb{R}^{p}$ are the dual variables for the inequality and equality constraints, respectively. This results in QP subproblems of the form:
$\operatorname*{minimize}_{dx}\quad$ $\frac{1}{2}dx^{\top}\nabla_{x}^{2}\mathcal{L}(x^{k},y_{I}^{k},y_{E}^{k})\,dx+\nabla f(x^{k})^{\top}dx,$ (2a)
$\operatorname*{subject\ to}\quad$ $g(x^{k})+\partial g(x^{k})dx\leq 0,\hskip 5.0pth(x^{k})+\partial h(x^{k})dx=0.$ (2b)
@@
Notably, the linearization of the constraints $g$ and $h$ may not produce a QP subproblem that is feasible, meaning that there may not exist any $dx$ that satisfies the linearized constraints Eq. 2b. In this case, the SQP solver either terminates with a suboptimal point or a specialized routine needs to be run in order to reduce the infeasibility. For example, when SNOPT encounters an infeasible subproblem, it enters elastic mode and solves a new optimization where the constraints are relaxed using $\ell_{1}$ penalty functions (Gill et al., 2005). This is advantageous over other choices, such as an $\ell_{2}$ penalty, as the $\ell_{1}$ norm encourages sparsity in the constraint violation. This means the optimizer can naturally identify the constraints that are the most difficult to satisfy. In the context of mixed integer NLPs, infeasibilities are very likely to occur during the branch and bound process, so their fast and robust identification is crucial (Gill and Wong, 2011).
@@
Moreover, as the interest in data-driven optimization grows, we often wish to solve many optimization problems with similar structure repeatedly (Amos and others, 2023), and potentially in parallel or in a batched fashion, such as in the method by Fang et al. (2023), which requires solving coupled systems of SQP problems in parallel. Needing to run specialized procedures in these cases is not scalable. Furthermore, identifying the best hyperparameters for each individual optimization problem is difficult and time consuming. Deep unfolding is a learning-to-optimize approach (Chen et al., 2022b) that has roots in the signal and image processing domains (Gregor and LeCun, 2010; Wang et al., 2015) and utilizes data-driven machine learning to reduce the cost of hyperparameter tuning and to accelerate convergence of model-based optimizers (Monga et al., 2021). It constitutes the state-of-the-art approach for sparse recovery (Liu et al., 2019) and video reconstruction (De Weerdt et al., 2024). In the context of NLP, deep unfolding has been recently applied to accelerate QPs. Saravanos et al. (2025) use an analogy to closed-loop control and learn feedback policies for the parameters of a deep-unfolded variant of the operator splitting QP (OSQP) solver (Stellato et al., 2020), which is a first-order method based on the alternating direction method of multipliers (ADMM) (Boyd et al., 2011). Their method can achieve orders-of-magnitude improvement in wall-clock time compared to OSQP, and they also propose a decentralized version for quickly solving QPs with distributed structure. Their idea is similar in vein to that of Ichnowski et al. (2021), who use reinforcement learning to train a policy that outputs the optimal parameters for OSQP, with the goal of accelerating the optimizer. Another related approach learns to warm-start a Douglas-Rachford splitting QP solver, with the goal of improving convergence speed (Sambharya et al., 2023).
Based on these considerations, in Section 3 we propose FlexQP, an always-feasible QP optimizer, and in Section 4 we apply deep unfolding, leading to Deep FlexQP. A more comprehensive survey of related work is provided in Appendix A.
@@ 3 FlexQP: An Always-Feasible Quadratic Programming Solver
Our proposed QP solver, FlexQP, transforms the QP constraints using an exact relaxation and then solves the resultant problem using an operator splitting inspired by OSQP (Stellato et al., 2020). We will assume the reader is familiar with ADMM; a good overview is provided by Boyd et al. (2011).
@@ 3.1 Quadratic Programming
We are interested in solving convex QPs of the general form
$\operatorname*{minimize}_{x}\hskip 5.0pt\frac{1}{2}x^{\top}Px+q^{\top}x,\quad\operatorname*{subject\ to}\hskip 5.0ptGx\leq h,\hskip 5.0ptAx=b,$ (3)
@@
The optimality conditions for Eq. 3 are given by:
$Px+q+G^{\top}y_{I}+A^{\top}y_{E}=0,\hskip 5.0ptGx-h\leq 0,\hskip 5.0ptAx-b=0,\hskip 5.0pty_{I}\odot(Gx-h)=0,\hskip 5.0pty_{I}\geq 0,$ (4)
where $y_{I}\in\mathbb{R}^{m}$ and $y_{E}\in\mathbb{R}^{p}$ are the dual variables for the inequality and equality constraints, respectively. A tuple $(x^{*},y_{I}^{*},y_{E}^{*})$ satisfying Eq. 4 is a solution to Eq. 3.
Throughout this work, we will avoid making any assumptions on $G$ or $A$, meaning that the constraints may be redundant, and in the worst case, there may not exist a feasible $x$ for the optimization. This allows for a unified way to handle infeasibilities when optimizations of the form Eq. 3 are embedded as a subproblem in SQP, as in Eq. 2.
@@ 3.2 Elastic Formulation
We start by introducing slack variables $s\in{\mathbb{R}}^{m}$, so Eq. 3 can be expressed equivalently as
$\operatorname*{minimize}_{x,s\geq 0}\hskip 5.0pt\frac{1}{2}x^{\top}Px+q^{\top}x,\quad\operatorname*{subject\ to}\hskip 5.0ptGx+s-h=0,\hskip 5.0ptAx-b=0.$ (5)
This standard technique is the basis of many interior point algorithms (Nocedal and Wright, 2006). We then relax the set of equality constraints in Eq. 5 using $\ell_{1}$ penalty functions, yielding
@@
with elastic penalty parameters $\mu_{I},\mu_{E}>0$. This relaxation approach is known as elastic programming (Brown and Graves, 1975), and one of the most well-known SQP-based solvers, SNOPT, uses this technique in order to reduce the infeasibility of a QP subproblem (Gill et al., 2005). This relaxation is also a fundamental step in the sequential $\ell_{1}$ quadratic programming method of Fletcher (1985). Notably, if Eq. 3 has a feasible solution and the elastic penalty parameters are sufficiently large, then the solutions to Eq. 3 and Eq. 6 are identical — this is why the relaxation is exact. On the other hand, if the original QP Eq. 3 is infeasible, then solving Eq. 6 finds a point that minimizes the constraint violation (Nocedal and Wright, 2006). This is formalized through the following theorem, which also describes what we mean by a sufficiently large penalty parameter.
@@
Let $(x^{*},y_{I}^{*},y_{E}^{*})$ solve Eq. 3. Let $\mu_{I}^{*}=\left\lVert y_{I}^{*}\right\rVert_{\infty}$ and $\mu_{E}^{*}=\left\lVert y_{E}^{*}\right\rVert_{\infty}$. Then, for all $\mu_{I}\geq\mu_{I}^{*}$ and $\mu_{E}\geq\mu_{E}^{*}$, the minimizers of Eq. 3 and Eq. 6 coincide.
This theorem is a generalization of the one by Han and Mangasarian (1979) that shows we can select a different penalty parameter for the inequality vs. equality constraints. The proof, provided in Appendix B, relies on two simple facts: the optimality conditions of Eq. 3 and the convexity of the objective. The proof also shows that it is possible to select vectors of penalty parameters $\mu_{I}$ and $\mu_{E}$, as long as each $\mu_{I,i}$ and $\mu_{E,j}$ obeys the constraints $\mu_{I,{i}}\geq|y_{I,i}|$ or $\mu_{E,j}\geq|y_{E,j}|$, respectively.
What happens when the penalty parameters $\bm{\mu}$ do not satisfy the conditions of Theorem 3.1? Using the interpretation of the Lagrange multiplier $y_{i}$ representing the cost of an individual constraint $i$ with associated penalty parameter $\mu_{i}>0$, if $\mu_{i}\geq|y_{i}|$ then the penalty on violating constraint $i$ in Eq. 6 is large enough such that Theorem 3.1 holds. On the other hand, if $\mu_{i}<|y_{i}|$, then this constraint $i$ is not being penalized strong enough, and so the solution to Eq. 6 will violate this constraint, with the amount of violation proportional to the difference between $\mu_{i}$ and $|y_{i}|$. We use this interpretation in Section 4 to design feedback policies that select the best penalty parameters as a function of the optimizer state and enforce the condition $\mu_{i}\geq|y_{i}|$ during learning using a supervised loss that includes the Lagrange multipliers (see also Theorem 3.3 below).